6x^3 = 18x + 60x^2
SOLVE

Answer by checkley71(8403)

6X^3=18X+60X^2
6X^3-60X^2-18X=0
6X(X^2-10X-3)=0
USING THE QUADRATIC EQUATION WE GET
X=(-B+-SQRT[B^2-4*A*C])/2A
X=(10+-SQRT[100-4*1*-3])/2*1
X=(10+-SQRT100+12)/2
X=(10+-SQRT112)/2
X=(10+-10.583)/2
X=(10+10.583)/2
X=20.583/2
X=10.2915 ANSWER
X=(10-10.583)/2
X=-.2915 ANSWER